### v1

@@ Abstract.
This paper presents IMAGGarment-1, a fine-grained garment generation (FGG) framework that enables high-fidelity garment synthesis with precise control over silhouette, color, and logo placement. Unlike existing methods that are limited to single-condition inputs, IMAGGarment-1 addresses the challenges of multi-conditional controllability in personalized fashion design and digital apparel applications. Specifically, IMAGGarment-1 employs a two-stage training strategy to separately model global appearance and local details, while enabling unified and controllable generation through end-to-end inference. In the first stage, we propose a global appearance model that jointly encodes silhouette and color using a mixed attention module and a color adapter. In the second stage, we present a local enhancement model with an adaptive appearance-aware module to inject user-defined logos and spatial constraints, enabling accurate placement and visual consistency. To support this task, we release GarmentBench, a large-scale dataset comprising over 180K garment samples paired with multi-level design conditions, including sketches, color references, logo placements, and textual prompts. Extensive experiments demonstrate that our method outperforms existing baselines, achieving superior structural stability, color fidelity, and local controllability performance. The code and model are available at https://github.com/muzishen/IMAGGarment-1.
@@
Fine-Grained Garment Generation; Multi-Conditional Generation; Fashion Design Applications; GarmentBench Dataset
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: October; June 03–05, 2025; Dublin, Irelandprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Visualization design and evaluation methods
Figure 1. Comparison of (a) existing sketch-to-image and logo insertion tasks with (b) our proposed fine-grained garment generation (FGG) task, which enables precise and controllable synthesis of garment structure, color, logo, and spatial placement. Unlike previous tasks that rely on a single input condition, FGG is tailored for real-world fashion design workflows by integrating multiple conditional controls.
1. Introduction
@@
Fine-Grained garment generation (FGG) aims to synthesize high-quality garments with precise control over garment silhouette, color scheme, logo content, and spatial placement. As personalized fashion and the digital apparel market grow rapidly, fine-grained controllability is increasingly crucial for applications in fashion design, brand customization, and personalized e-commerce.
@@
Recently, image synthesis (Podell et al., 2023; Lipman et al., 2022) has made notable progress in tasks such as sketch-to-image generation (Voynov et al., 2023; Isola et al., 2017) and logo insertion (Kim et al., 2024) (as illustrated in Figure 1 (a)), demonstrating basic capabilities in structural and content-level control. However, these tasks provide only coarse guidance and rely on single-condition inputs (e.g., sketch or color), lacking the fine-grained controllability needed to model the nuanced interactions between global structure and local details in garment design. Although sequential or modular combinations may offer partial solutions, they fail to explicitly disentangle and jointly model global attributes (e.g., silhouette, color) and local appearance details (e.g., logo content and spatial placement). Without unified control mechanisms or hierarchical representations, these approaches often suffer from condition entanglement, conflicting objectives, and visual inconsistencies, ultimately falling short of the high standards required in real-world fashion design. In contrast, practical fashion design requires joint control over multiple interdependent factors: designers determine global attributes such as silhouette and color, followed by fine-tuning of local elements like logos and their placement. To support this process, a unified generation task that clearly separates and coordinates global and local attributes is essential for controllable and high-fidelity synthesis.
@@
To address these limitations, we propose a new task: fine-grained garment generation (FGG), as illustrated in Figure 1 (b). FGG is formulated as a unified multi-conditional garment synthesis task, taking a textual prompt, garment silhouette, color palette, and spatially constrained logos as joint inputs. It aims to generate garments that faithfully reflect high-level structural intent and fine-grained local styling cues. FGG is specifically designed to mirror real-world fashion workflows, where designers must coordinate diverse input modalities to express creative intent. Unlike conventional approaches that process each condition independently or sequentially, FGG emphasizes joint modeling and hierarchical reasoning across input types. It goes beyond simple task combinations by enforcing consistent integration of global and local attributes within a unified generation framework, enabling nuanced control over the overall structure and detailed appearance. Specifically, FGG task introduces three key challenges: (1) maintaining visual and semantic consistency across heterogeneous input conditions, (2) resolving conflicts between global structures and localized visual elements, and (3) generalizing to unseen condition combinations without retraining (see Figure 1(c)). FGG thus marks a fundamental shift from single-condition or loosely coupled pipelines toward a unified, design-intent-driven generation paradigm that better reflects the complexity of real-world garment design.
@@
To this end, we propose IMAGGarment-1, a two-stage training and end-to-end inference framework tailored for fine-grained garment generation. Unlike prior methods (Ye et al., 2023; Zhang et al., 2023a; Rombach et al., 2022) that rely on single-condition inputs or simple feature fusion, our framework is explicitly designed to achieve fine-grained controllability under multiple, interdependent constraints. In the first stage, we propose a global appearance model with a mixed attention module and a color adapter to jointly encode garment silhouette and color palette, improving overall appearance fidelity and mitigating condition entanglement. In the second stage, we present a local enhancement model equipped with an adaptive appearance-aware module to inject user-defined logos and their spatial constraints, enabling precise logo placement while preserving global consistency. To further promote research in this direction, we release GarmentBench, a large-scale dataset comprising over 180k garment samples annotated with rich multi-level design conditions, including silhouette sketches, color references, logo placements, and textual prompts. Extensive experiments demonstrate that IMAGGarment-1 significantly outperforms existing baselines in terms of structural stability and local controllability. To summarize, the main contributions are listed as follows:
@@
  • We propose IMAGGarment-1, a controllable garment generation framework that enables precise control over garment structure, color, and logo placement, addressing the challenges of FGG.
@@
  • We release GarmentBench, a large-scale dataset with diverse garments and rich multi-conditional annotations, serving as a valuable benchmark for FGG research.
2. Related Work
2.1. GAN-Based Methods
@@
Early approaches (Chen et al., 2020, 2021; Wu et al., 2022; Ghosh et al., 2019; Chen and Hays, 2018; Ghosh et al., 2019; Li et al., 2020) to garment generation often employ generative adversarial networks (GANs) (Creswell et al., 2018; Mirza and Osindero, 2014; Men et al., 2020), focusing primarily on sketch-to-image translation (Liu et al., 2017) by learning spatial mappings from structural cues. For example, DeepFaceDrawing (Chen et al., 2020) and DeepFaceEditing (Chen et al., 2021) decompose sketches into individual facial components and progressively assemble them into photorealistic portraits. Then, DeepPortraitDrawing (Wu et al., 2022) extends this strategy to full-body synthesis using a local-to-global pipeline. Similarly, Interactive frameworks (Ghosh et al., 2019) introduce gating mechanisms to support user-driven editing, while DALColor (Ci et al., 2018) combines WGAN-GP with line-art colorization for refined appearance control. However, these methods (Chen et al., 2020, 2021; Wu et al., 2022; Ghosh et al., 2019) are limited to single-condition generation and are difficult to apply in real-world fashion scenarios that demand multi-conditional controllability. Moreover, GAN-based approaches (Ma et al., 2017; Men et al., 2020; Siarohin et al., 2018) are prone to adversarial training instability and image artifacts, often compromising realism.
2.2. Diffusion-Based Methods
Diffusion models (Ho et al., 2020; Song et al., 2020; Gao et al., 2025) have made significant strides in conditional image generation, thanks to their iterative denoising mechanism and flexible control capacity. To enhance controllability, plugin-based methods such as IP-Adapter (Ye et al., 2023), ControlNet (Zhang et al., 2023a), and BLIP-Diffusion (Li et al., 2023) inject external conditions into pre-trained diffusion models via lightweight modules. Parallel architectures (Shen et al., 2024; Lin et al., 2024; Chen et al., 2024a; Xu et al., 2024) further improve control precision by incorporating reference-guided features during generation. In the fashion domain, DiffCloth (Zhang et al., 2023b) enables localized garment editing via part-specific textual prompts, allowing independent control over different regions (e.g., sleeves, collars). For logo generation, AnyLogo (Zhang et al., 2024) employs a dual-state denoising strategy to preserve fine logo details; LogoSticker (Zhu et al., 2024) introduces token-based injection for flexible logo placement; and RefDiffuser (Kim et al., 2024) leverages expert-driven plugins to enhance texture and spatial fidelity. However, these methods typically focus on either global appearance or localized editing in isolation, and lack a unified framework for jointly modeling multiple design conditions in fine-grained garment synthesis.
3. Methodology
@@ 3.1. Overall Framework
To address the above challenges, we propose IMAGGarment-1, a conditional diffusion framework tailored for fine-grained garment generation. Our framework comprises two components: a global appearance model (stage I) and a local enhancement model (stage II), which explicitly disentangle and jointly control the global appearance and local details under multi-conditional guidance, enabling accurate synthesis of garment silhouette, color, and logo placement. As illustrated in Figure 2, the global appearance model first generates a latent of coarse garment image conditioned on the textual prompt, garment silhouette, and color palette. Subsequently, the local enhancement model refines this latent representation by integrating user-defined logo and spatial constraint, producing the final high-fidelity garment image with fine-grained controllability. Specifically, the global appearance model (Section 3.2) leverages our proposed mixed attention module and color adapter to effectively capture global appearance features from textual descriptions, silhouettes, and colors, while mitigating entanglement among these conditions. The local enhancement model (Section 3.3) introduces an adaptive appearance-aware module ($A^{3}$ Module) that injects logo content and spatial location constraint into the latent space, achieving precise logo placement. Finally, the training and inference strategies used in IMAGGarment-1 are summarized in Section 3.4.
@@
Figure 2. Visualization of the IMAGGarment-1 inference pipeline. The global appearance model generates coarse latent from textual prompts, silhouettes, and colors. The local enhancement model then injects user-defined logos and spatial location constraints to produce the fine-grained garment.
3.2. Stage I: Global Appearance Model
@@
Figure 3. Overview of our IMAGGarment-1 framework. IMAGGarment-1 is a two-stage conditional diffusion framework for fine-grained garment generation. The global appearance model first synthesizes a coarse latent representation from the input text prompt, silhouette, and color palette using a parallel UNet with mixed attention and a color adapter. The local enhancement model then refines this latent by injecting user-defined logos and location constraints through the proposed $A^{3}$ module, enabling precise logo placement and high-fidelity garment generation.
@@
Motivation. Existing garment generation methods (Ye et al., 2023; Zhang et al., 2023a; Rombach et al., 2022) typically rely on single-condition inputs (e.g., sketch or text), causing entangled features and limited controllability. To resolve this, we propose a global appearance model that explicitly disentangles silhouette, color, and text, enabling precise multi-conditional control.
@@
Architecture. As shown in the left of the Figure 3, our global appearance model comprises two shared frozen VAE encoders, one frozen VAE decoder, a trainable silhouette UNet, a frozen text encoder, a trainable color adapter, and a denoising UNet with the proposed mixed attention. Specifically, we first utilize the frozen VAE encoder to project the input reference silhouette into the latent space. Subsequently, we employ a trainable silhouette UNet (structurally identical to the denoising UNet but without cross attention) to extract fine-grained silhouette features, which are then integrated into the frozen denoising UNet via our proposed mixed attention module. Meanwhile, textual features obtained from the frozen CLIP text encoder and color features extracted by the proposed color adapter are further fused into the denoising UNet through cross attention. After multiple denoising iterations, the model generates coarse garment images that precisely align with the reference silhouette and faithfully reflect user-specified color.
@@
Mixed Attention. To effectively incorporate reference silhouette features into the denoising UNet without compromising the generative capability of the original UNet, we propose a mixed attention module. As shown in Figure 3, we extend all self attention layers in the denoising UNet to the proposed mixed attention, which introduces two additional learnable projection layers to align the silhouette features $C_{s}$ with the latent features $Z_{t}$. Formally, the mixed attention is defined as:
@@
(1) $Z_{m}=\text{Softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)V+\alpha\cdot\text{Softmax}\left(\frac{Q(K^{\prime})^{T}}{\sqrt{d}}\right)V^{\prime},$
@@
(2) $Q=Z_{t}W_{q},\ K=Z_{t}W_{k},\ V=Z_{t}W_{v},\ K^{\prime}=C_{s}W^{\prime}_{k},\ V^{\prime}=C_{s}W^{\prime}_{v}$
@@
(3) $Z_{new}=\text{Softmax}\left(\frac{QK_{t}^{T}}{\sqrt{d}}\right)V_{t}+\beta\cdot\text{Softmax}\left(\frac{QK_{c}^{T}}{\sqrt{d}}\right)V_{c},\vspace{-0.2cm}$
@@
3.3. Stage II: Local Enhancement Model
Motivation. Existing methods (Zhang et al., 2024; Zhu et al., 2024) typically neglect detailed logo integration or treat it as a separate task, causing poor spatial alignment and visual inconsistency. To address this limitation, we propose a local enhancement model equipped with an adaptive appearance-aware ($A^{3}$) module, explicitly injecting user-defined logos and spatial constraints into the latent space. This design enables precise, consistent control over localized garment details, significantly enhancing visual fidelity.
Architecture. As illustrated on the right of Figure 3, the local enhancement model comprises a frozen VAE encoder and decoder, a denoising UNet, and an adaptive appearance-aware module ($A^{3}$ module). The $A^{3}$ module fuses local conditions, such as logos and spatial constraints, by concatenating them along spatial or channel dimensions, enabling precise control over fine-grained visual elements. Given a garment, logo, and placement mask, the model adaptively adjusts the logo’s size and position while preserving its visual fidelity. To reduce redundancy and focus on local detail refinement, we optimize only the self attention layers of the denoising UNet and discard all cross attention layers, as the global appearance model has already encoded the textual information.

### v2

@@ Abstract
This paper presents IMAGGarment, a fine-grained garment generation (FGG) framework that enables high-fidelity garment synthesis with precise control over silhouette, color, and logo placement. Unlike existing methods that are limited to single-condition inputs, IMAGGarment addresses the challenges of multi-conditional controllability in personalized fashion design and digital apparel applications. Specifically, IMAGGarment employs a two-stage training strategy to separately model global appearance and local details, while enabling unified and controllable generation through end-to-end inference. In the first stage, we propose a global appearance model that jointly encodes silhouette and color using a mixed attention module and a color adapter. In the second stage, we present a local enhancement model with an adaptive appearance-aware module to inject user-defined logos and spatial constraints, enabling accurate placement and visual consistency. To support this task, we release GarmentBench, a large-scale dataset comprising over 180K garment samples paired with multi-level design conditions, including sketches, color references, logo placements, and textual prompts. Extensive experiments demonstrate that our method outperforms existing baselines, achieving superior structural stability, color fidelity, and local controllability performance. Code, models, and datasets are publicly available at https://github.com/muzishen/IMAGGarment.
Index Terms:
@@
Fine-Grained Garment Generation, Multi-Conditional Generation, Fashion Design Applications, GarmentBench Dataset.
I Introduction
@@
Fine-Grained garment generation (FGG) aims to synthesize high-quality garments with precise control over garment silhouette, color scheme, logo content, and spatial placement. As personalized fashion and the digital apparel market grow rapidly, fine-grained controllability [1, 2, 3, 4] is increasingly crucial for applications in fashion design and e-commerce.
In traditional garment ideation [5, 6] and visualization [7, 8], designers analyze line drawings to establish silhouette and construction, then select color palettes and materials, and finally arrange brand elements such as logos and trims. This manual workflow has two persistent drawbacks. First, it is time consuming: to match the specification, edits must be applied object by object and view by view; in a seasonal collection, even identical panels within the same board are recolored or relabeled one at a time, which does not scale. Second, it is error prone and inconsistent: small deviations in hue, shading, or logo placement arise across artists and rounds of revision, yielding mismatches across styles, sizes, and camera viewpoints. As project scope grows, these issues inflate turnaround time and complicate quality control and version management.
@@
Recently, image synthesis [9, 10, 11, 12] has made notable progress in tasks such as sketch-to-image generation [13, 14, 15, 16] and logo insertion [17, 18, 19] (as illustrated in Fig. 1 (a)), demonstrating basic capabilities in structural and content-level control. However, these tasks [13, 17, 20] provide only coarse guidance and rely on single-condition inputs (e.g., sketch or color), lacking the fine-grained controllability needed to model the nuanced interactions between global structure and local details in garment design. Although sequential or modular combinations may offer partial solutions, they [21, 22, 23] fail to explicitly disentangle and jointly model global attributes (e.g., silhouette, color) and local appearance details (e.g., logo content and spatial placement). Without unified control mechanisms, these approaches [21, 22, 23] often suffer from condition entanglement, conflicting objectives, and visual inconsistencies, ultimately falling short of the high standards required in real-world fashion design. In contrast, practical fashion design [5, 6] requires joint control over multiple interdependent factors: designers determine global attributes such as silhouette and color, followed by fine-tuning of local elements like logos and their placement. To support this process, a unified generation task that clearly separates and coordinates global and local attributes is essential for controllable and high-fidelity synthesis.
Figure 1: Comparison of (a) existing sketch-to-image and logo insertion tasks with (b) our proposed fine-grained garment generation (FGG) task, which enables precise and controllable synthesis of garment structure, color, logo, and spatial placement. Unlike previous tasks that rely on a single input condition, FGG is tailored for real-world fashion design workflows by integrating multiple conditional controls.
@@
To address these limitations, we propose a new task: fine-grained garment generation (FGG), as illustrated in Fig. 1 (b). FGG is formulated as a unified multi-conditional garment synthesis task, taking a textual prompt, garment silhouette, color palette, and spatially constrained logos as joint inputs. It aims to generate garments that faithfully reflect high-level structural intent and fine-grained local styling cues. FGG is specifically designed to mirror real-world fashion workflows, where designers must coordinate diverse input modalities to express creative intent. Unlike conventional approaches that process each condition independently or sequentially, FGG emphasizes joint modeling and hierarchical reasoning across input types. It goes beyond simple task combinations by enforcing consistent integration of global and local attributes within a unified generation framework, enabling nuanced control over the overall structure and detailed appearance. Specifically, FGG task introduces three key challenges: (1) maintaining visual and semantic consistency across heterogeneous input conditions, (2) resolving conflicts between global structures and localized visual elements, and (3) generalizing to unseen condition combinations without retraining (see Fig. 1(c)). FGG thus marks a fundamental shift from single-condition or loosely coupled pipelines toward a unified, design-intent-driven generation paradigm that better reflects the complexity of real-world garment design.
@@
To this end, we propose IMAGGarment, a two-stage training and end-to-end inference framework tailored for fine-grained garment generation. Unlike prior methods that rely on single-condition inputs or simple feature fusion, our framework is explicitly designed to achieve fine-grained controllability under multiple, interdependent constraints. In the first stage, we propose a global appearance model with a mixed attention module and a color adapter to jointly encode garment silhouette and color palette, improving overall appearance fidelity and mitigating condition entanglement. In the second stage, we present a local enhancement model equipped with an adaptive appearance-aware module to inject user-defined logos and their spatial constraints, enabling precise logo placement while preserving global consistency. To further promote research in this direction, we release GarmentBench, a large-scale dataset comprising over 180k garment samples annotated with rich multi-level design conditions, including silhouette sketches, color references, logo placements, and textual prompts. Extensive experiments demonstrate that IMAGGarment significantly outperforms existing baselines in terms of structural stability and local controllability. To summarize, the main contributions are listed as follows:
@@
  • We propose IMAGGarment, a controllable garment generation framework that enables precise control over garment structure, color, and logo placement, addressing the challenges of FGG.
@@
  • We release GarmentBench, a large-scale dataset with diverse garments and rich multi-conditional annotations, serving as a valuable benchmark for controllable garment generation research.
The remainder of this paper is organized as follows. Section II surveys prior work on garment generation, encompassing GAN-based techniques and diffusion-based controllable generation. Section III describes the proposed IMAGGarment methodology, comprising a global appearance model with mixed attention and a color adapter, a local enhancement model with the A3 module, and the associated training and inference strategies. Section IV presents the experimental protocol and results, including the GarmentBench dataset and evaluation metrics, implementation details, and results and analysis. Section V concludes the paper.
II Related Work
II-A GAN-Based Methods
@@
Early approaches [24, 25, 26, 27, 28, 29] to garment generation predominantly build on generative adversarial networks (GANs)[30, 31, 32], with a major line devoted to sketch-to-image translation[33] that learns spatial mappings from structural cues. Representative systems such as DeepFaceDrawing [24] and DeepFaceEditing [25] decompose sketches into semantic components and progressively assemble photorealistic results, while DeepPortraitDrawing [26] extends this paradigm to full-body synthesis via local-to-global pipelines. Interactive frameworks [27] further introduce gating mechanisms for user-guided editing, and DALColor [34] combines WGAN-GP [35] with line-art colorization for refined appearance control. Beyond sketches, related GAN-based efforts explore pose- or part-guided generation [36, 37], leveraging learned warping or deformable alignment to better propagate structural constraints from sources to targets.
However, these methods [24, 25, 26, 27] are largely restricted to single-condition settings (e.g., sketches or poses alone), making it difficult to support real-world fashion scenarios that require joint control over multiple factors such as silhouette, garment layers, color/pattern, and local embellishments. Moreover, adversarial training is prone to instability and visual artifacts [36, 32, 37], and the reliance on paired or carefully aligned supervision limits robustness to occlusion, diverse body shapes, and open-world catalogs. As a result, while GAN-based pipelines can produce plausible textures under constrained conditions, they struggle to achieve reliable, fine-grained, and multi-conditional controllability at scale.
II-B Diffusion-Based Methods
Diffusion models [38, 39, 40] have achieved strong progress in conditional image generation owing to their iterative denoising process and flexible conditioning interfaces. To improve controllability with minimal modification to large backbones, plugin-based approaches such as IP-Adapter [21], ControlNet [22], and BLIP-Diffusion [41] inject external conditions (e.g., reference images, structural maps, or language cues) through lightweight adapters. In parallel, reference-guided or dual-stream designs [42, 43, 44, 45] propagate features from exemplars alongside text/image prompts, thereby strengthening identity preservation and fine control during sampling.
In fashion-related applications, DiffCloth [46] supports localized garment edits via part-specific textual prompts, enabling independent control over regions such as sleeves and collars. For logo-centric generation, AnyLogo [18] adopts a dual-state denoising strategy to retain subtle logo details; LogoSticker [19] performs token-based injection to flexibly place logo elements; and RefDiffuser [17] leverages expert-driven plugins to enhance texture fidelity and spatial alignment. Despite these advances, most methods emphasize either global appearance control or localized editing in isolation. A unified framework that jointly models multiple design conditions, e.g., silhouette and layer topology together with color/pattern and local embellishments, while maintaining structural coherence across the denoising trajectory remains underexplored.
TABLE I: Definitions of main symbols used in this paper.
Notation Definition
t{t} Timestep
$Z_{t}$} Latent feature at $t$ step
$Z_{m}$} Output of mixed attention
$x_{0}$} Real image
$x_{t}$} Noisy data at $t$ step
G{G} Garment image
L{L} Logo image
M{M} Mask image
$C_{g}$} Feature of garment image
$C_{l}$} Feature of logo image
$C_{m}$} Feature of mask image
$C_{s}$} Feature of silhouette image
$C_{c}$} Feature of color image
$C_{t}$} Feature of text prompt
$\theta_{g}$} Global appearance model
$\theta_{l}$} Local enhancement model
$\epsilon$} Gaussian noise
$\bar{\alpha}_{t}$} Cumulative product of noise weights
w{w} Guidance scale
$\alpha$} Silhouette scale
$\beta$} Color scale
III Methodology
Symbol Definition. To introduce our IMAGGarment method more clearly, we define the main symbols used throughout the paper in TABLE I.
@@ III-A Overall Framework
To address the above challenges, we propose IMAGGarment, a conditional diffusion framework tailored for fine-grained garment generation. Our framework comprises two components: a global appearance model (stage I) and a local enhancement model (stage II), which explicitly disentangle and jointly control the global appearance and local details under multi-conditional guidance, enabling accurate synthesis of garment silhouette, color, and logo placement. As illustrated in Fig. 2, the global appearance model first generates a latent of coarse garment image conditioned on the textual prompt, garment silhouette, and color palette. Subsequently, the local enhancement model refines this latent representation by integrating user-defined logo and spatial constraint, producing the final high-fidelity garment image with fine-grained controllability. Specifically, the global appearance model (Section III-B) leverages our proposed mixed attention module and color adapter to effectively capture global appearance features from textual descriptions, silhouettes, and colors, while mitigating entanglement among these conditions. The local enhancement model (Section III-C) introduces an adaptive appearance-aware module ($A^{3}$ Module) that injects logo content and spatial location constraint into the latent space, achieving precise logo placement. Finally, the training and inference strategies used in IMAGGarment are summarized in Section III-D.
@@
Figure 2: Visualization of the IMAGGarment inference pipeline. The global appearance model generates coarse latent from textual prompts, silhouettes, and colors. The local enhancement model then injects user-defined logos and spatial location constraints to produce the fine-grained garment.
III-B Stage I: Global Appearance Model
@@
Figure 3: Overview of our IMAGGarment framework. IMAGGarment is a two-stage conditional diffusion framework for fine-grained garment generation. The global appearance model first synthesizes a coarse latent representation from the input text prompt, silhouette, and color palette using a parallel UNet with mixed attention and a color adapter. The local enhancement model then refines this latent by injecting user-defined logos and location constraints through the proposed $A^{3}$ module, enabling precise logo placement and high-fidelity garment generation.
@@
Motivation. Existing garment generation methods [21, 22, 23] typically rely on single-condition inputs (e.g., sketch or text), causing entangled features and limited controllability. To resolve this, we propose a global appearance model that explicitly disentangles silhouette, color, and text, enabling precise multi-conditional control.
@@
Architecture. As shown in the left of the Fig. 3, our global appearance model comprises two shared frozen VAE encoders, one frozen VAE decoder, a trainable silhouette UNet, a frozen text encoder, a trainable color adapter, and a denoising UNet with the proposed mixed attention. Specifically, we first utilize the frozen VAE encoder to project the input reference silhouette into the latent space. Subsequently, we employ a trainable silhouette UNet (structurally identical to the denoising UNet but without cross attention) to extract fine-grained silhouette features, which are then integrated into the frozen denoising UNet via our proposed mixed attention module. Meanwhile, textual features obtained from the frozen CLIP text encoder and color features extracted by the proposed color adapter are further fused into the denoising UNet through cross attention. After multiple denoising iterations, the model generates coarse garment images that precisely align with the reference silhouette and faithfully reflect user-specified color.
@@
Mixed Attention. To effectively incorporate reference silhouette features into the denoising UNet without compromising the generative capability of the original UNet, we propose a mixed attention module. As shown in Fig. 3, we extend all self attention layers in the denoising UNet to the proposed mixed attention, which introduces two additional learnable projection layers to align the silhouette features $C_{s}$ with the latent features $Z_{t}$. Formally, the mixed attention is defined as:
@@
$Z_{m}=\text{Softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)V+\alpha\cdot\text{Softmax}\left(\frac{Q(K^{\prime})^{T}}{\sqrt{d}}\right)V^{\prime},$ (1)
@@
$Q=Z_{t}W_{q},\ K=Z_{t}W_{k},\ V=Z_{t}W_{v},\ K^{\prime}=C_{s}W^{\prime}_{k},\ V^{\prime}=C_{s}W^{\prime}_{v}$ (2)
@@
$Z_{new}=\text{Softmax}\left(\frac{QK_{t}^{T}}{\sqrt{d}}\right)V_{t}+\beta\cdot\text{Softmax}\left(\frac{QK_{c}^{T}}{\sqrt{d}}\right)V_{c},$ (3)
@@
III-C Stage II: Local Enhancement Model
Motivation. Existing methods [18, 19] typically neglect detailed logo integration or treat it as a separate task, causing poor spatial alignment and visual inconsistency. To address this limitation, we propose a local enhancement model equipped with an adaptive appearance-aware ($A^{3}$) module, explicitly injecting user-defined logos and spatial constraints into the latent space. This design enables precise, consistent control over localized garment details, significantly enhancing visual fidelity.
Architecture. As illustrated on the right of Fig. 3, the local enhancement model comprises a frozen VAE encoder and decoder, a denoising UNet, and an adaptive appearance-aware module ($A^{3}$ module). The $A^{3}$ module fuses local conditions, such as logos and spatial constraints, by concatenating them along spatial or channel dimensions, enabling precise control over fine-grained visual elements. Given a garment, logo, and placement mask, the model adaptively adjusts the logo’s size and position while preserving its visual fidelity. To reduce redundancy and focus on local detail refinement, we optimize only the self attention layers of the denoising UNet and discard all cross attention layers, as the global appearance model has already encoded the textual information.
Figure 4: Overview of GarmentBench dataset construction pipeline and samples. (a) Data construction pipeline for GarmentBench. (b) Example samples with multimodal annotations: silhouette, logo, text, logo location, and color.